# Parametric Expensive Multi-Objective Optimization via Generative Solution Modeling

### v2

@@ B.1 Synthetic Benchmarks
We evaluate our approach on three classical multi-objective test problems from the DTLZ suite: DTLZ-1, DTLZ-2, and DTLZ-3 [DTLZ]. These scalable benchmark functions are widely used for evaluating multi-objective optimization algorithms due to their well-understood theoretical properties and configurable Pareto Front geometries [DTLZ].
To create parametric families of these problems, we introduce the task parameter $\boldsymbol{\theta}\in[0.8,1]$. The task parameter $\boldsymbol{\theta}$ controls the power scaling applied to the decision variables $\mathbf{x}$ in the objective function calculations, creating meaningful variations in the problem landscape through power transformations of the form $F(\mathbf{x}^{\boldsymbol{\theta}})$. This parameterization enables the exploration of related optimization instances that share a similar structure but exhibit different convergence characteristics and Pareto Front properties.
@@ B.2 Real-World Benchmarks
The design variables $\mathbf{x}$ encode the geometric structure of a three-arm symmetric desk lamp, including base dimensions, arm templates, and hand orientations that are rotated to create the complete fixture. The operational condition $\boldsymbol{\theta}$ represents the desired focal point height, spanning from desk-level to floor-lamp configurations across a continuous range. We seek to balance three potentially conflicting objectives: structural stability (minimizing wobbliness under perturbations), material efficiency (reducing overall weight and cost), and illumination quality (achieving optimal light coverage at the target location).
@@
The design variables $\mathbf{x}$ control the elevation profile of interior grid points on a rooftop solar installation, allowing for curved or stepped surface geometries while maintaining fixed boundary conditions. The operational condition $\boldsymbol{\theta}$ captures the orientation of the building relative to the cardinal directions in a continuous rotational space. The optimization balances energy harvest during morning hours (capturing early sunlight for peak demand periods) against evening energy collection (storing power for nighttime usage), while maintaining structural smoothness for practical construction and aesthetic appeal.
@@
The design variables $\mathbf{x}$ amount to thrust commands for a twin-rotor UAV across multiple time steps, defining the complete flight trajectory from initial state to target destination. The operational conditions $\boldsymbol{\theta}$ involve the vehicle’s physical characteristics, including fuselage length and structural density, representing different UAV configurations. The controller must simultaneously minimize navigation accuracy (how close the vehicle gets to its target destination) and efficiency (conserving battery power for extended flight duration), representing the classic trade-off between precision and endurance in aerial robotics [bicopter].
@@
The design variables $\mathbf{x}$ determine the geometric parameters of a biomedical microdevice, including channel gap width, magnetic element length, and structural thickness. The operational conditions $\boldsymbol{\theta}$ represent the magnetic properties of both cell types that need to be separated, varying across a continuous range of real values. The device must achieve three conflicting objectives: purity (minimizing capture of unwanted cell types), selectivity (maximizing the separation quality between desired and undesired cells), and manufacturing efficiency (reducing device thickness to minimize fabrication cost) [extremo]. This represents a practical biomedical engineering challenge where the same device design must be optimized for separating different patient samples or cell types with varying magnetic characteristics from complex biological mixtures such as blood.
Appendix C Experimental Details
@@ C.1 Generative Model Training
For the DDPM, we use a simple MLP architecture that takes the concatenated input $\mathbf{x},t,\boldsymbol{c}]$, where $t$ is the timestep, and $\mathbf{c}$ is the conditioning vector. The network consists of 4 fully connected layers with hidden dimension 128, following the structure of [Diffuse-BBO]. We employ a linear noise schedule with $\beta$ values ranging from 0.0001 to 0.02 over 1000 time steps. The model is trained to predict the noise $\boldsymbol{\epsilon}$ added during the forward diffusion process using standard MSE loss. During training, we randomly sample timesteps $t\sim\text{Uniform}(0,T-1)$ and compute the loss $\mathcal{L}=\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\phi}(\mathbf{x}_{t},t,\mathbf{c})\|_{2}^{2}$. For sampling, we use the standard DDPM reverse process, starting from Gaussian noise and iteratively denoising over all timesteps. We apply gradient clipping with a maximum norm of 1.0 for training stability and clamp the generated samples to the $[0,1]$ range to handle boundary constraints.
We employed Adam optimizer [kingma2017adammethodstochasticoptimization] for both models and the learning rate is set to 0.1 for VAE and 0.001 for DDPM.
@@ C.2 GP Model Training
For the Gaussian Process model, we implement a composite kernel GP that decomposes the input space into decision variables $\mathbf{x}\in\mathbb{R}^{D}$ and task parameters $\boldsymbol{\theta}\in\Theta\subset\mathbb{R}^{V}$. The model uses exact GP inference with a Gaussian likelihood, and the related hyperparameters are trained using the exact marginal log-likelihood optimized by Adam [kingma2017adammethodstochasticoptimization].
@@
where $\kappa_{\text{dec}}$ is an isotropic RBF kernel for decision variables with lengthscale constrained to the interval $[0.1,2.5]$, and $\kappa_{\text{task}}$ is an ARD RBF kernel for task parameters, allowing each task dimension to have its own lengthscale parameter. The overall covariance function is wrapped by a scale kernel to scale the product of the two kernels. This composite kernel is partially inspired by the kernel function in [offline-contextual]. The GP-based source codes are partially supported by GPytorch [Gpytorch].

### v3

@@ Appendix B Benchmark Definitions
In Table 3, we specify the dimension of solution and task spaces for each benchmark problem as well as the number of objectives. Note that, since the problem formulation we focus on is expensive multi-objective optimization problems, too large a solution or task space might easily incur a dimensionality curse for those surrogate models and generative models. Therefore, in this paper, we mainly focus on solving low-to-moderate P-EMOPs. And we leave solving high-dimensional P-EMOPs as a future challenging direction here.
@@ B.1 Synthetic Benchmarks
We evaluate our approach on three classical multi-objective test problems from the DTLZ suite: DTLZ-1, DTLZ-2, and DTLZ-3 [13]. These scalable benchmark functions are widely used for evaluating multi-objective optimization algorithms due to their well-understood theoretical properties and configurable Pareto Front geometries [13].
To create parametric families of these problems, we introduce the task parameter $\boldsymbol{\theta}\in[0.8,1]$. The task parameter $\boldsymbol{\theta}$ controls the power scaling applied to the decision variables $\mathbf{x}$ in the objective function calculations, creating meaningful variations in the problem landscape through power transformations of the form $F(\mathbf{x}^{\boldsymbol{\theta}})$. This parameterization enables the exploration of related optimization instances that share a similar structure but exhibit different convergence characteristics and Pareto Front properties. To compare the multi-objective optimization performance, the reference point for each benchmark is set to $[200.0,200.0]$, $[2.0,2.0]$, and $[240.0,240.0]$.
@@ B.2 Real-World Benchmarks
The design variables $\mathbf{x}$ encode the geometric structure of a three-arm symmetric desk lamp, including base dimensions, arm templates, and hand orientations that are rotated to create the complete fixture. The operational condition $\boldsymbol{\theta}$ represents the desired focal point height, spanning from desk-level to floor-lamp configurations across a continuous range. We seek to balance three potentially conflicting objectives: structural stability (minimizing wobbliness under perturbations), material efficiency (reducing overall weight and cost), and illumination quality (achieving optimal light coverage at the target location). To compare the multi-objective optimization performance, the reference point for this problem is set to $[1.0,1.0,1.0]$.
@@
The design variables $\mathbf{x}$ control the elevation profile of interior grid points on a rooftop solar installation, allowing for curved or stepped surface geometries while maintaining fixed boundary conditions. The operational condition $\boldsymbol{\theta}$ captures the orientation of the building relative to the cardinal directions in a continuous rotational space. The optimization balances energy harvest during morning hours (capturing early sunlight for peak demand periods) against evening energy collection (storing power for nighttime usage), while maintaining structural smoothness for practical construction and aesthetic appeal. To compare the multi-objective optimization performance, the reference point for this problem is set to $[1.0,1.0]$.
@@
The design variables $\mathbf{x}$ amount to thrust commands for a twin-rotor UAV across multiple time steps, defining the complete flight trajectory from initial state to target destination. The operational conditions $\boldsymbol{\theta}$ involve the vehicle’s physical characteristics, including fuselage length and structural density, representing different UAV configurations. The controller must simultaneously minimize navigation accuracy (how close the vehicle gets to its target destination) and efficiency (conserving battery power for extended flight duration), representing the classic trade-off between precision and endurance in aerial robotics [52]. To compare the multi-objective optimization performance, the reference point for this problem is set to $[1.0,1.0]$.
@@
The design variables $\mathbf{x}$ determine the geometric parameters of a biomedical microdevice, including channel gap width, magnetic element length, and structural thickness. The operational conditions $\boldsymbol{\theta}$ represent the magnetic properties of both cell types that need to be separated, varying across a continuous range of real values. The device must achieve three conflicting objectives: purity (minimizing capture of unwanted cell types), selectivity (maximizing the separation quality between desired and undesired cells), and manufacturing efficiency (reducing device thickness to minimize fabrication cost) [33]. This represents a practical biomedical engineering challenge where the same device design must be optimized for separating different patient samples or cell types with varying magnetic characteristics from complex biological mixtures such as blood. To compare the multi-objective optimization performance, the reference point for this problem is set to $[1.0,1.0,1.0]$.
Appendix C Experimental Setup
@@ C.1 Generative Model Training
For the DDPM, we use a simple MLP architecture that takes the concatenated input $\mathbf{x},t,\boldsymbol{c}]$, where $t$ is the timestep, and $\mathbf{c}$ is the conditioning vector. The network consists of 4 fully connected layers with hidden dimension 128, following the structure of [28]. We employ a linear noise schedule with $\beta$ values ranging from 0.0001 to 0.02 over 1000 time steps. The model is trained to predict the noise $\boldsymbol{\epsilon}$ added during the forward diffusion process using standard MSE loss. During training, we randomly sample timesteps $t\sim\text{Uniform}(0,T-1)$ and compute the loss $\mathcal{L}=\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\phi}(\mathbf{x}_{t},t,\mathbf{c})\|_{2}^{2}$. For sampling, we use the standard DDPM reverse process, starting from Gaussian noise and iteratively denoising over all timesteps. We apply gradient clipping with a maximum norm of 1.0 for training stability and clamp the generated samples to the $[0,1]$ range to handle boundary constraints.
We employed Adam optimizer [26] for both models and the learning rate is set to 0.1 for VAE and 0.001 for DDPM.
@@ C.2 GP Model Training
For the Gaussian Process model, we implement a composite kernel GP that decomposes the input space into decision variables $\mathbf{x}\in\mathbb{R}^{D}$ and task parameters $\boldsymbol{\theta}\in\Theta\subset\mathbb{R}^{V}$. The model uses exact GP inference with a Gaussian likelihood, and the related hyperparameters are trained using the exact marginal log-likelihood optimized by Adam [26].
@@
where $\kappa_{\text{dec}}$ is an isotropic RBF kernel for decision variables with lengthscale constrained to the interval $[0.1,2.5]$, and $\kappa_{\text{task}}$ is an ARD RBF kernel for task parameters, allowing each task dimension to have its own lengthscale parameter. The overall covariance function is wrapped by a scale kernel to scale the product of the two kernels. This composite kernel is partially inspired by the kernel function in [6]. The GP-based source codes are partially supported by GPytorch [16].
Appendix D Detailed Experimental Results
D.1 Ablation Studies
Variants DTLZ-1 DTLZ-2 DTLZ-3
PMT-MOBO (w/o GENERATIVE) 1.709e+04 (4.696e+03) 2.767e+00 (5.500e-02) 3.785e+03 (5.397e+03)
PLAIN-VAE (w/o ACQUISITION) 2.052e+04 (7.555e+03) 2.933e+00 (2.315e-01) 1.301e+04 (8.679e+03)
PMT-MOBO-VAE 3.101e+04 (3.415e+03) 3.052e+00 (9.580e-02) 2.098e+04 (5.335e+03)
Table 4: Ablation Study of VAE Integration. The Alternating Loop (PMT-MOBO-VAE) consistently outperforms both pure Acquisition Search and pure Generative Sampling.
Variants DTLZ-1 DTLZ-2 DTLZ-3
PMT-MOBO (w/o GENERATIVE) 1.709e+04 (4.696e+03) 2.767e+00 (5.500e-02) 3.785e+03 (5.397e+03)
PLAIN-DDPM (w/o ACQUISITION) 1.384e+04 (3.847e+03) 3.052e+00 (3.970e-02) 2.337e+03 (3.425e+03)
PMT-MOBO-DDPM 3.374e+04 (8.354e+02) 3.131e+00 (1.620e-02) 3.093e+04 (1.673e+03)
Table 5: Ablation Study of DDPM Integration. The proposed PMT-MOBO-DDPM consistently achieves superior performance, particularly on deceptive landscapes (DTLZ-3) where plain generation fails.
Tables 4 and 5 present the ablation results for VAE and DDPM integrations, respectively. We isolate the impact of the alternating loop by comparing the full method against two decoupled baselines: pure acquisition search (PMT-MOBO w/o GENERATIVE) and pure generative sampling (PLAIN-VAE/DDPM). The results for the VAE variant (Table 4) clearly illustrate the limitations of isolated components. PLAIN-VAE exhibits a high standard deviation across benchmarks (e.g., $7.56\text{e}3$ on DTLZ-1), indicating that while pure generative sampling can occasionally "get lucky" and hit high-quality solutions, it suffers from inherent stochasticity during the optimization process and lacks the precision to consistently refine the Pareto front. Conversely, pure acquisition search often converges prematurely to local optima, failing to explore promising but unseen regions of the design space. By coupling these components, PMT-MOBO-VAE leverages generative sampling to enhance candidate diversity while utilizing the alternating acquisition search to lower the variance of pure sampling. This synergy results in statistically superior convergence and stability across all tested problems.
Variants ($Q\%$) DTLZ-1 DTLZ-2 DTLZ-3
PMT-MOBO-VAE-S (5%) 2.523e+04 (4.644e+03) 3.046e+00 (6.060e-02) 2.256e+04 (7.200e+03)
PMT-MOBO-VAE (10%) 3.101e+04 (3.415e+03) 3.052e+00 (9.580e-02) 2.098e+04 (5.335e+03)
PMT-MOBO-VAE-L (20%) 2.802e+04 (3.211e+03) 2.995e+00 (7.380e-02) 1.576e+04 (9.457e+03)
Table 6: Sensitivity Analysis of Elite Ratio ($Q\%$) for VAE. The base setting (10%) is generally most robust, although PMT0-MOBO-VAE-S (5%) excels on DTLZ-3 by avoiding local optimum overfitting but resulting in a larger variance.
Variants ($Q\%$) DTLZ-1 DTLZ-2 DTLZ-3
PMT-MOBO-DDPM-S (5%) 3.362e+04 (7.397e+02) 3.091e+00 (1.630e-02) 2.721e+04 (1.790e+03)
PMT-MOBO-DDPM (10%) 3.374e+04 (8.354e+02) 3.131e+00 (1.620e-02) 3.093e+04 (1.673e+03)
PMT-MOBO-DDPM-L (20%) 3.367e+04 (6.563e+02) 3.099e+00 (1.000e-02) 2.888e+04 (2.341e+03)
Table 7: Sensitivity Analysis of Elite Ratio ($Q\%$) for DDPM. The proposed method demonstrates high stability, with the default configuration (10%) consistently yielding peak Hypervolume.
D.2 Sensitivity Analysis
We investigate the sensitivity of the elite ratio parameter $Q$, which governs the trade-off between exploration breadth and exploitation intensity.This trend is most pronounced in the VAE sensitivity analysis shown in Table 6.
  • Large $Q$ (20%): Increasing the elite ratio dilutes the quality of the training set. This reduces the selective pressure, causing the generative model to learn the distribution of sub-optimal, low-quality solutions rather than focusing on the high-performing elite regions. Consequently, performance drops on DTLZ-1 and DTLZ-3.
  • Small $Q$ (5%): Conversely, an overly strict selection criterion limits the training data availability. While this theoretically sharpens exploitation, in practice, it prevents the model from learning a well-behaved latent structure, incurring large variance and instability (e.g., higher standard deviation on DTLZ-1).
  • Optimal $Q$ (10%): The default setting strikes an effective balance, providing sufficient data for stable training while maintaining high selective pressure to guide the model toward the Pareto front.
While the DDPM variant (Table 7) shows greater resilience to parameter changes, the underlying trade-off remains consistent: $Q=10\%$ consistently yields the most robust hypervolume performance.
Component PMT-MOBO-VAE (s) PMT-MOBO-DDPM (s)
GP Training $7.92\pm 2.64$ $7.19\pm 2.25$
Acquisition Optimization $6.10\pm 0.34$ $5.90\pm 0.31$
Generative Model Training $0.55\pm 0.26$ $0.39\pm 0.15$
Generative Solution Sampling $6.08\pm 0.28$ $6.03\pm 0.50$
Table 8: Runtime Overhead Analysis on DTLZ-2 ($T=50$). The table reports the wall-clock time (Mean $\pm$ Std) for each algorithmic component per active iteration. Acquisition Optimization occurs only in ACQUISITION mode, while generative model training and solution sampling occur only in GENERATIVE mode. GP Training occurs in every iteration.
D.3 Computational Overhead
A potential concern with integrating deep generative models into P-EMOPs is the computational cost of iterative retraining. To address this, we profiled the wall-clock time of each algorithmic component on the DTLZ-2 benchmark over $T=50$ iterations. As detailed in Table 8, the cost of generative model training is negligible (<0.6<0.6s per iteration), orders of magnitude faster than the standard Gaussian Process (GP) training ($\approx 7.9$s). Furthermore, the generative solution sampling ($\approx 6.0$s) incurs a computational cost comparable to standard acquisition optimization ($\approx 6.1$s). Since our alternating framework swaps between acquisition and generative modes rather than running them simultaneously, the total runtime per iteration remains consistent with standard MOBO methods. This confirms that our framework enhances optimization performance without imposing a prohibitive computational burden.